# Chat2SVG: Vector Graphics Generation with Large Language Models and Image Diffusion Models

### v1

@@ Abstract
Scalable Vector Graphics (SVG) has become the de facto standard for vector graphics in digital design, offering resolution independence and precise control over individual elements. Despite their advantages, creating high-quality SVG content remains challenging, as it demands technical expertise with professional editing software and a considerable time investment to craft complex shapes. Recent text-to-SVG generation methods aim to make vector graphics creation more accessible, but they still encounter limitations in shape regularity, generalization ability, and expressiveness. To address these challenges, we introduce Chat2SVG, a hybrid framework that combines the strengths of Large Language Models (LLMs) and image diffusion models for text-to-SVG generation. Our approach first uses an LLM to generate semantically meaningful SVG templates from basic geometric primitives. Guided by image diffusion models, a dual-stage optimization pipeline refines paths in latent space and adjusts point coordinates to enhance geometric complexity. Extensive experiments show that Chat2SVG outperforms existing methods in visual fidelity, path regularity, and semantic alignment. Additionally, our system enables intuitive editing through natural language instructions, making professional vector graphics creation accessible to all users. The project page is https://chat2svg.github.io.
@@ 1 Introduction
Existing approaches have explored image-based methods for text-to-SVG generation, iteratively optimizing a large collection of shape elements (e.g., cubic Bézier curves, typically to ) by rendering them into images with differentiable rasterizers [20] and evaluating them using text-image similarity metrics like CLIP loss [28] and Score Distillation Sampling (SDS) loss [27]. While these image-based methods [9, 16, 43, 44, 50] can generate visually impressive SVG through the combination of numerous paths and powerful image models, they face a critical limitation in maintaining the regularity and semantics of shapes. Specifically, semantic components that should be represented as single elements often end up fragmented across multiple overlapping paths. Although these fragmented paths, when viewed collectively, can yield visually appealing outputs, they fundamentally conflict with professional design principles, where each semantic component is intentionally crafted as a single, regularized path.
@@
Given SVG is defined using Extensible Markup Language (XML), another solution to text-to-SVG synthesis involves language-based methods, which treat SVG scripts as text input. Recent works [41, 35] have proposed specialized tokenization strategies to embed SVG scripts and trained Sequence-To-Sequence (seq2seq) models on domain-specific datasets (e.g., icons and fonts). While these approaches achieve good generation quality within their training domains, they suffer from limited generalization due to the absence of large-scale, general-purpose text-SVG training data. Meanwhile, some works [26, 31] show that while state-of-the-art Large Language Models (LLMs) can generate basic geometric shapes (e.g., circles and rectangles) with layouts matching text prompts, they struggle to produce the complex geometric details required for professional SVG applications. Consequently, the limited generalization and poor expressiveness of language-based methods hinder their adoption for text-to-SVG generation.
@@
To address the aforementioned problems, we propose a novel hybrid framework that leverages the complementary strengths of image-based and language-based methods. Our approach first utilizes an LLM to synthesize SVG templates composed of basic primitives and then optimizes their geometric details guided by image diffusion models. This hybrid framework addresses the key limitations of both paradigms: The LLM-based template generation overcomes the domain-specific generalization constraints of traditional language models while ensuring shape regularity, as each element naturally corresponds to a single semantic component. The subsequent image-based optimization then enhances the expressiveness of these well-structured templates by capturing complex geometric details that LLMs alone struggle to produce. Specifically, to fully exploit LLMs’ ability to create reasonable and complex layouts that match text prompts, we design an SVG-oriented prompt with multiple stages, including prompt expansion, SVG script generation, and visual refinement. To refine the generated SVG templates, we implement a dual-stage optimization strategy: (1) We first use an image diffusion model to synthesize detailed image-level targets, then follow [50] to optimize path-level latent vectors with a pretrained SVG VAE, eliminating common issues such as self-intersections and jagged curves; (2) We further refine geometric details by directly optimizing the point coordinates to capture fine-grained visual elements. Comprehensive experiments show that Chat2SVG consistently outperforms existing methods regarding overall visual quality, individual path regularity, and semantic coherence. Furthermore, our system enables intuitive editing through iterative natural language instructions, making vector graphics creation more accessible to non-expert users. Our contributions are summarized as follows:
@@ 2.1 Large Language Models for Design
LLMs have emerged as powerful tools for graphics design tasks, demonstrating remarkable capabilities in understanding and reasoning about complex design specifications across diverse domains. Recent works have explored their applications in (1) visual layout and composition, including poster designs [34, 5, 47, 21], 3D scene arrangements [38, 14, 33, 51, 32, 2, 10], and placement of basic shapes for image generation [8, 46]; (2) content creation and manipulation for shapes [11], materials [15], and animation [22, 36]; and (3) design understanding [18, 42]. While these demonstrate LLMs’ versatility, their potential for vector graphics remains unexplored. Our work addresses this gap by showing that LLMs can effectively generate structured SVG templates, extending their capabilities to the domain of vector graphics creation.
@@ 2.2 Text-Guided Vector Graphics Generation
Text-to-SVG generation approaches can be roughly categorized into image-based and language-based methods. Image-based methods [9, 16, 43, 44, 50] start with a large collection of randomly initialized shapes, render vector graphics with a differentiable renderer [20], and iteratively optimize path parameters by minimizing text-image similarity losses (e.g., CLIP [28] loss and Score Distillation Sampling [27] loss). Despite their visually pleasant appearance, these methods often generate SVG results containing multiple fragmented paths that lack individual semantic correspondence, which is undesirable for real design scenarios. Language-based methods [41, 3, 35] design specialized tokenization strategies to encode vector graphics into discrete tokens, which are then concatenated with text tokens, and train seq2seq models like autoregressive transformers [37] on domain-specific datasets (e.g., icons and fonts). While these approaches are conceptually elegant, their generalization ability is limited by the scarcity of large-scale, general-purpose vector graphics datasets. Some works [26, 31] have explored LLMs’ vector graphics generation capabilities, showing that while LLMs can create reasonable layouts matching text prompts, they struggle to produce complex geometric shapes. Our Chat2SVG combines image-based and language-based methods to address the limitations of each paradigm, producing semantically meaningful SVG where each path corresponds to a distinct visual element and contains refined geometric details.
@@ 2.3 Vector Graphics Representation Learning
Vector graphics representation learning is essential for sketch-based retrieval, reconstruction, and generation tasks. A seminal study, SketchRNN [12], combines a Recurrent Neural Network (RNN) and a VAE to learn vector representations. Lopes et al. [24] trained a VAE to represent image-level font styles, then passed the latent embedding into a decoder for vector font generation. To increase reconstruction quality, later approaches like Sketchformer [30] and DeepSVG [4] adopted Transformer [37] architectures, followed by dual-modality methods [39, 23, 40] that leveraged both vector and image features. However, encoding the entire complex SVG into a single latent embedding often causes detail loss. To address this limitation, Zhang et al. [50] proposed path-level representational learning, creating a smooth latent space for individual paths. We leverage these path-level latent embeddings in our optimization to eliminate self-intersecting and jagged paths.
@@ 3.1 SVG Template Generation with LLMs
Prompt Expansion. Users often provide vague and brief prompts to the LLM when describing their desired graphics. Such unstructured input may lead to oversimplified SVG (regarding both the number of elements and structures) that fails to reflect the intended design. Consequently, we propose a three-layer prompt expansion strategy that systematically delineates the initial prompts: (1) Scene-level: Starting with an initial prompt, we instruct the LLM to analyze it holistically and identify essential objects that should appear. The LLM then expands the prompt by suggesting complementary objects to enhance scene completeness. (2) Object-level: For each object in the expanded prompt, we guide the LLM to systematically break it down into its components. For instance, when describing a lion, the LLM deconstructs it into distinct parts such as the body, head, mane, legs, eyes, ears, and tail. This decomposition ensures that no critical components are missing during SVG script generation. (3) Layout-level: After object decomposition, we direct the LLM to develop a comprehensive layout plan. This includes determining the position and size of each element on the canvas, selecting appropriate colors for visual harmony, and establishing clear spatial relationships to ensure a cohesive overall composition.
SVG Script Generation. After prompt expansion, we obtain a detailed scene description. We then convert this natural language specification into SVG scripts in XML format. Since the LLM has limitations in synthesizing geometrically complex paths, we constrain the shapes to a carefully selected set of basic primitives: rectangles, ellipses, lines, polylines, polygons, and short paths. We fix the canvas size to . Each geometric element is assigned a unique ID and includes semantic annotations in its comments.
@@ 3.2 SVG Optimization Guided by Image Diffusion
Detail Enhancement. While the primitive shapes in the SVG template have accurate semantic meaning, they lack the sophisticated geometric details necessary for a professional design. To strengthen visual expressiveness, we first render the SVG template into a raster image. We then use an image editing method to generate a more detailed version that serves as our optimization target. Specifically, we employ SDEdit [25], which enhances the input image by first adding noise and then progressively denoising it with an image diffusion model to produce outputs with richer details, such as the contour of the lion’s mane illustrated in Figure 2. To maintain structural similarity between the enhanced output and the original image, we incorporate a ControlNet (tile version) [49] into our pipeline. By using Gaussian-blurred versions of the initial images as control signals, the ControlNet effectively maintains the overall composition throughout the enhancement process. Apart from geometric details, this image-to-image translation process also introduces new semantic parts such as the lion’s beard. To faithfully reproduce the target image, we propose using the Segment Anything Model (SAM) [17] to identify and add these new parts into the SVG template. In particular, we denote the rendered images of the initial SVG template and its diffusion-enhanced version as and , respectively. We apply SAM to both images to obtain their corresponding mask sets: for the template image and for the target image, where and represent the number of masks in each set. The major idea of path addition is that the diffusion process primarily introduces small decorative details (e.g., the beard and toes of the lion in Figure 2) while maintaining the key components (e.g., the tail and the body). Therefore, we implement a mask-filtering strategy to identify these new elements: For each mask in , we compare it against all masks in using Intersection over Union (IoU) loss [29]. If overlaps with a template mask and their IoU is below a threshold, we identify as a new decorative detail that should be added to the SVG. Conversely, if their IoU exceeds the threshold, we interpret as a shape variation of an existing component and exclude it from further processing. Finally, for all filtered masks, we approximate their boundaries with polygons and add them to the SVG script.
@@
Notation. We define an SVG script as a collection of paths (i.e., shapes), . Each path consists of a sequence of commands, . A command is defined by its type and its associated control points , where M represents Move and C represents cubic Bézier curves. For consistency, we convert all other primitive shapes (e.g., rectangles and ellipses) in the SVG template into cubic Bézier curves.
@@
Previous image-based methods [9, 16, 43, 44] optimize SVG shapes by directly manipulating the control points . Despite convenient, this approach often leads to artifacts such as self-intersecting curves and unnatural deformations. To overcome these limitations, Zhang et al. [50] introduced a path-level SVG VAE. This model’s latent space effectively captures common shape patterns and geometric constraints, allowing for the optimization of path latent vectors at a higher level to produce smooth outputs.
@@
Latent Optimization. We leverage this pretrained SVG VAE to conduct the latent optimization. However, their SVG encoder expects a fixed number of commands ( cubic Bézier curves), while our SVG primitives contain varying, typically fewer commands. To resolve this mismatch, we develop a latent inversion process that converts our primitives into latent embeddings. We start with a randomly sampled latent vector and decode it into a command sequence. Then we evenly sample points along its contour. Similarly, we sample an equal number of points along the contour of our target primitive shape. Using these point sets, we compute the Earth Mover’s Distance (EMD) [1, 7, 48]:
@@
After obtaining latent vectors for all paths, we jointly optimize these vectors along with visual attribute parameters: filling colors , stroke properties (color and width) , and transformation matrices . At each optimization iteration , we decode the latent vectors into command sequences and apply transformation matrices to the control points. We then combine color and stroke attributes together, and use a differentiable rasterizer [20] to render the complete SVG into an image . Our optimization objective consists of three loss terms: (1) An MSE loss between the rendered image and target image to ensure visual similarity; (2) A curvature loss that reduces sharp bends and fluctuations along the contour by computing the discrete second derivative:
@@
Point Optimization. While latent optimization effectively aligns SVG shapes with their target positions and contours, it has a notable limitation: the resulting shapes tend to be overly smooth, lacking the intricate details that are often essential in professional vector graphics. For example, in Figure 2, the lion’s mane after latent optimization does not capture the fine contour present in the target image. Therefore, we add a second optimization stage that directly refines the control points of the paths, similar to previous image-based works [9, 16, 43, 44]. To achieve finer granularity in shape control, we split each cubic Bézier curve at its midpoint, effectively doubling the number of control points in the SVG. This increased control point density allows for more precise shape adjustments. The point optimization stage employs a loss function that combines the MSE loss and the curvature loss:

### v2

@@ Abstract
Scalable Vector Graphics (SVG) has become the de facto standard for vector graphics in digital design, offering resolution independence and precise control over individual elements. Despite their advantages, creating high-quality SVG content remains challenging, as it demands technical expertise with professional editing software and a considerable time investment to craft complex shapes. Recent text-to-SVG generation methods aim to make vector graphics creation more accessible, but they still encounter limitations in shape regularity, generalization ability, and expressiveness. To address these challenges, we introduce Chat2SVG, a hybrid framework that combines the strengths of Large Language Models (LLMs) and image diffusion models for text-to-SVG generation. Our approach first uses an LLM to generate semantically meaningful SVG templates from basic geometric primitives. Guided by image diffusion models, a dual-stage optimization pipeline refines paths in latent space and adjusts point coordinates to enhance geometric complexity. Extensive experiments show that Chat2SVG outperforms existing methods in visual fidelity, path regularity, and semantic alignment. Additionally, our system enables intuitive editing through natural language instructions, making professional vector graphics creation accessible to all users. Our code is available at https://chat2svg.github.io/.
@@ 1 Introduction
Existing approaches have explored image-based methods for text-to-SVG generation, iteratively optimizing a large collection of shape elements (e.g., cubic Bézier curves, typically to ) by rendering them into images with differentiable rasterizers [Li2020DiffVG] and evaluating them using text-image similarity metrics like CLIP loss [radford2021learning] and Score Distillation Sampling (SDS) loss [poole2022dreamfusion]. While these image-based methods [frans2021clipdraw, jain2022vectorfusion, xing2023diffsketcher, xing2024svgdreamer, zhang2024text] can generate visually impressive SVG through the combination of numerous paths and powerful image models, they face a critical limitation in maintaining the regularity and semantics of shapes. Specifically, semantic components that should be represented as single elements often end up fragmented across multiple overlapping paths. Although these fragmented paths, when viewed collectively, can yield visually appealing outputs, they fundamentally conflict with professional design principles, where each semantic component is intentionally crafted as a single, regularized path.
@@
Given SVG is defined using Extensible Markup Language (XML), another solution to text-to-SVG synthesis involves language-based methods, which treat SVG scripts as text input. Recent works [wu2023iconshop, tang2024strokenuwa] have proposed specialized tokenization strategies to embed SVG scripts and trained Sequence-To-Sequence (seq2seq) models on domain-specific datasets (e.g., icons and fonts). While these approaches achieve good generation quality within their training domains, they suffer from limited generalization due to the absence of large-scale, general-purpose text-SVG training data. Meanwhile, some works [openai2023gpt4, sharma2024vision] show that while state-of-the-art Large Language Models (LLMs) can generate basic geometric shapes (e.g., circles and rectangles) with layouts matching text prompts, they struggle to produce the complex geometric details required for professional SVG applications. Consequently, the limited generalization and poor expressiveness of language-based methods hinder their adoption for text-to-SVG generation.
@@
To address the aforementioned problems, we propose a novel hybrid framework that leverages the complementary strengths of image-based and language-based methods. Our approach first utilizes an LLM to synthesize SVG templates composed of basic primitives and then optimizes their geometric details guided by image diffusion models. This hybrid framework addresses the key limitations of both paradigms: The LLM-based template generation overcomes the domain-specific generalization constraints of traditional language models while ensuring shape regularity, as each element naturally corresponds to a single semantic component. The subsequent image-based optimization then enhances the expressiveness of these well-structured templates by capturing complex geometric details that LLMs alone struggle to produce. Specifically, to fully exploit LLMs’ ability to create reasonable and complex layouts that match text prompts, we design an SVG-oriented prompt with multiple stages, including prompt expansion, SVG script generation, and visual refinement. To refine the generated SVG templates, we implement a dual-stage optimization strategy: (1) We first use an image diffusion model to synthesize detailed image-level targets, then follow [zhang2024text] to optimize path-level latent vectors with a pretrained SVG VAE, eliminating common issues such as self-intersections and jagged curves; (2) We further refine geometric details by directly optimizing the point coordinates to capture fine-grained visual elements. Comprehensive experiments show that Chat2SVG consistently outperforms existing methods regarding overall visual quality, individual path regularity, and semantic coherence. Furthermore, our system enables intuitive editing through iterative natural language instructions, making vector graphics creation more accessible to non-expert users. Our contributions are summarized as follows:
@@ 2.1 Large Language Models for Design
LLMs have emerged as powerful tools for graphics design tasks, demonstrating remarkable capabilities in understanding and reasoning about complex design specifications across diverse domains. Recent works have explored their applications in (1) visual layout and composition, including poster designs [tang2023layoutnuwa, cheng2024graphic, yang2024posterllava, lin2024layoutprompter], 3D scene arrangements [wang2024chat2layout, hu2024scenecraft, tam2024scenemotifcoder, zhang2023scenewiz3d, sun20233d, aguina2024open, fu2025anyhome], and placement of basic shapes for image generation [feng2024layoutgpt, yamada2024l3go]; (2) content creation and manipulation for shapes [ganeshan2024parsel], materials [huang2024blenderalchemy], and animation [liu2024logomotion, tseng2024keyframer]; and (3) design understanding [kulits2024re, wu2024gpt]. While these demonstrate LLMs’ versatility, their potential for vector graphics remains unexplored. Our work addresses this gap by showing that LLMs can effectively generate structured SVG templates, extending their capabilities to the domain of vector graphics creation.
@@ 2.2 Text-Guided Vector Graphics Generation
Text-to-SVG generation approaches can be roughly categorized into image-based and language-based methods. Image-based methods [frans2021clipdraw, jain2022vectorfusion, xing2023diffsketcher, xing2024svgdreamer, zhang2024text] start with a large collection of randomly initialized shapes, render vector graphics with a differentiable renderer [Li2020DiffVG], and iteratively optimize path parameters by minimizing text-image similarity losses (e.g., CLIP [radford2021learning] loss and Score Distillation Sampling [poole2022dreamfusion] loss). Despite their visually pleasant appearance, these methods often generate SVG results containing multiple fragmented paths that lack individual semantic correspondence, which is undesirable for real design scenarios. Language-based methods [wu2023iconshop, belouadi2023automatikz, tang2024strokenuwa] design specialized tokenization strategies to encode vector graphics into discrete tokens, which are then concatenated with text tokens, and train seq2seq models like autoregressive transformers [vaswani2017attention] on domain-specific datasets (e.g., icons and fonts). Although these approaches are conceptually elegant, their generalizability is limited by the scarcity of large-scale, general-purpose vector graphics datasets. Some works [openai2023gpt4, sharma2024vision] have explored LLMs’ vector graphics generation capabilities, showing that while LLMs can create reasonable layouts matching text prompts, they struggle to produce complex geometric shapes. Our Chat2SVG combines image-based and language-based methods to address the limitations of each paradigm, producing semantically meaningful SVG where each path corresponds to a distinct visual element and contains refined geometric details.
@@ 2.3 Vector Graphics Representation Learning
Vector graphics representation learning is essential for sketch-based retrieval, reconstruction, and generation tasks. A seminal study, SketchRNN [ha2017neural], combines a Recurrent Neural Network (RNN) and a VAE to learn vector representations. lopes2019learned trained a VAE to represent image-level font styles, then passed the latent embedding into a decoder for vector font generation. To increase reconstruction quality, later approaches like Sketchformer [ribeiro2020sketchformer] and DeepSVG [carlier2020deepsvg] adopted Transformer [vaswani2017attention] architectures, followed by dual-modality methods [wang2021deepvecfont, liu2023dualvector, wang2023deepvecfont] that leveraged both vector and image features. However, encoding the entire complex SVG into a single latent embedding often causes detail loss. To address this limitation, zhang2024text proposed path-level representational learning, creating a smooth latent space for individual paths. We leverage these path-level latent embeddings in our optimization to eliminate self-intersecting and jagged paths.
@@ 3.1 SVG Template Generation with LLMs
Prompt Expansion. Users often provide vague and brief prompts to the LLM when describing their desired graphics. Such unstructured input may lead to oversimplified SVG (regarding both the number of elements and structures) that fails to reflect the intended design. Consequently, we propose a three-level prompt expansion strategy that systematically delineates the initial prompts: (1) Scene-level: Starting with an initial prompt, we instruct the LLM to analyze it holistically and identify essential objects that should appear. The LLM then expands the prompt by suggesting complementary objects to enhance scene completeness. (2) Object-level: For each object in the expanded prompt, we guide the LLM to systematically break it down into its components. For instance, when describing a lion, the LLM deconstructs it into distinct parts such as the body, head, mane, legs, eyes, ears, and tail. This decomposition ensures that no critical components are missing during SVG script generation. (3) Layout-level: After object decomposition, we direct the LLM to develop a comprehensive layout plan. This includes determining the position and size of each element on the canvas, selecting appropriate colors for visual harmony, and establishing clear spatial relationships to ensure a cohesive overall composition.
SVG Script Generation. After prompt expansion, we obtain a detailed scene description. We then convert this natural language specification into SVG scripts in XML format. Since the LLM has limitations in synthesizing geometrically complex paths, we constrain the shapes to a carefully selected set of basic primitives: rectangles, ellipses, lines, polylines, polygons, and short paths. We set the canvas size to . Each geometric element is assigned a unique ID and includes semantic annotations in its comments.
@@ 3.2 SVG Optimization Guided by Image Diffusion
Detail Enhancement. While the primitive shapes in the SVG template have accurate semantic meaning, they lack the geometric details necessary for a professional design. To improve visual expressiveness, we first render the SVG template into a raster image, and then use an image editing method to generate a more detailed version that serves as our optimization target. Specifically, we employ SDEdit [meng2021sdedit], which enhances the input image by first adding noise and then progressively denoising it with an image diffusion model to produce outputs with richer details, such as the contour of the lion’s mane illustrated in Figure 2. To maintain structural similarity between the enhanced output and the original image, we incorporate a ControlNet (tile version) [zhang2023adding] into our pipeline. By using Gaussian-blurred versions of the initial images as control signals, the ControlNet effectively maintains the overall composition throughout the enhancement process.
Apart from geometric details, this image-to-image translation process also introduces new semantic parts such as the lion’s beard. To faithfully reproduce the target image, we propose using the Segment Anything Model (SAM) [kirillov2023segment] to identify these new decorative details. Specifically, we denote the rendered images of the initial SVG template and its diffusion-enhanced counterpart as and , respectively. We apply SAM to extract mask sets from both images: for the template image and for the target image, where and represent the number of masks in each collection. Our approach is based on the observation that the diffusion process simultaneously generates fine decorative details that are absent in the original template (e.g., the beard of the lion in Figure 2) while preserving the structural integrity of larger components (such as the body). Consequently, we implement an area-based mask-filtering strategy to identify these newly generated elements. For each mask detected in , we evaluate it against all masks in using Intersection over Union (IoU) metrics [rezatofighi2019generalized]. When overlaps with a template mask but their IoU falls below a threshold, we classify as a new decorative component. Alternatively, if their IoU exceeds the threshold, we interpret as a variation of an existing component. For all filtered masks classified as new decorative shapes, we approximate their boundaries with polygons and integrate them into the SVG script.
@@
We define an SVG script as a collection of paths (i.e., shapes), . Each path consists of a sequence of commands, . A command is defined by its type and its associated control points , where M represents Move and C represents cubic Bézier curves. For consistency, we convert all other primitive shapes (e.g., rectangles and ellipses) in the SVG template into cubic Bézier curves.
@@
Previous image-based methods [frans2021clipdraw, jain2022vectorfusion, xing2023diffsketcher, xing2024svgdreamer] optimize SVG shapes by directly manipulating the control points . Despite convenience, this approach often leads to artifacts such as self-intersecting curves and unnatural deformations. To overcome these limitations, zhang2024text introduced a path-level SVG VAE. This model’s latent space effectively captures common shape patterns and geometric constraints, allowing for the optimization of path latent vectors at a higher level to produce smooth outputs.
@@
Latent Optimization. We leverage this pretrained SVG VAE to conduct the latent optimization. However, their SVG encoder expects a fixed number of commands ( cubic Bézier curves), while our SVG primitives contain varying (typically fewer) commands. To resolve this mismatch, we develop a latent inversion process that converts our primitives into latent embeddings. We start with a randomly sampled latent vector and decode it into a command sequence. Then we evenly sample points along its contour. Similarly, we sample an equal number of points along the contour of our target primitive shape. Using these point sets, we compute the Earth Mover’s Distance (EMD) [achlioptas2018learning, fan2017point, yu2018pu]:
@@
After obtaining latent vectors for all paths, we jointly optimize these vectors along with visual attribute parameters: filling colors , stroke properties (color and width) , and transformation matrices . At each optimization iteration , we decode the latent vectors into command sequences and apply transformation matrices to the control points. We then combine color and stroke attributes together, and use a differentiable rasterizer [Li2020DiffVG] to render the complete SVG into an image . Our optimization objective consists of three loss terms: (1) An MSE loss between the rendered image and target image to ensure visual similarity; (2) A curvature loss that reduces sharp bends and fluctuations along the contour by computing the discrete second derivative:
@@
Point Optimization. While latent optimization effectively aligns SVG shapes with their target positions and contours, it has a notable limitation: the resulting shapes tend to be overly smooth, lacking the intricate details that are often essential in professional vector graphics. For example, in Figure 2, the lion’s mane after latent optimization does not capture the fine contour present in the target image. Therefore, we add a second optimization stage that directly refines the control points of the paths, similar to previous image-based works [frans2021clipdraw, jain2022vectorfusion, xing2023diffsketcher, xing2024svgdreamer]. To achieve finer granularity in shape control, we split each cubic Bézier curve at its midpoint, effectively doubling the number of control points in the SVG. This increased control point density allows for more precise shape adjustments. The point optimization stage employs a loss function that combines the MSE loss and the curvature loss: